OpenStudy (anonymous):

An open rectangular box with volume 2m^3 has a square base. Express the surface area of the box as a function of the length of a side of the base.

OpenStudy (anonymous):

the answer the book gave was S(x)= x^2 + (8/x), x>0 but I don't understand this

OpenStudy (anonymous):

Start with the base of the box. You know its area is length x width, and that since its a square, the length of the each side of the base are equal to each other. So you can say length x length = the are of the base. Now each side of the box has the area length x height. Notice that "length" here means the same as the "length" on the base, because they are connected to the base :D. Think about this, try drawing a picture and make a formula for it. The last bit is just considering how it relates to the 2m^3 volume you've been given, but I think that depends on the level of maths you are in.

OpenStudy (anonymous):

calculus 1

OpenStudy (anonymous):

but how do you find surface area?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Sorry, working on my own homework. Im not sure I understand your question. The area of a 2d object is its surface area.

OpenStudy (mathstudent55):

Can you express the volume of the box using the figure above?

OpenStudy (mathstudent55):

@mangoes20 Are you still there?

OpenStudy (anonymous):

i'm just not understanding at alll

OpenStudy (mathstudent55):

Do you see my figure above?

OpenStudy (mathstudent55):

If you want help with this problem, I'll help you now. If you don't answer, I'm getting off in a few minutes.

OpenStudy (anonymous):

ok i'm ready

OpenStudy (mathstudent55):

OK, Look at the figure above. That is a box with a base and 4 sides but no top.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The surface area of the box is the sum of the areas of all 4 vertical faces and the area of the base.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The base of the box is a square with side length x. What is the area of the base?

OpenStudy (anonymous):

x x y

OpenStudy (mathstudent55):

No. That is the volume of the box which we'll need soon. The base of the box is a square. The area of a square is the square of the side. Since the side of the base has length x, the area of the base is x^2.

OpenStudy (mathstudent55):

Now we have the area of the base. The next thing we need is the area of the 4 vertical sides. Each vertical side is a rectangle with side lengths x and y. What is the area of each vertical side?

OpenStudy (anonymous):

xy^2

OpenStudy (mathstudent55):

How do you get xy^2 from multiplying x and y? Each side is a rectangle x by y, so its area is simply xy.

OpenStudy (mathstudent55):

There are 4 vertical sides, so the total are of vertical sides is 4xy.

OpenStudy (anonymous):

oh

OpenStudy (mathstudent55):

Now we have the area of the base is x^2, and the area of all 4 sides is 4xy. A = x^2 + 4xy

OpenStudy (mathstudent55):

We need the area expressed in terms of x only, but we have y in the expression. Now we use the volume to get rid of y.

OpenStudy (mathstudent55):

The volume of a box is V = LWH ( volume = length * width * height)

OpenStudy (mathstudent55):

For this box, we are given the volume is 2 m^3. The volume is V = x * x* y V = x^2y

OpenStudy (mathstudent55):

We are given that the volume is 2 m^2, so we write x^2y = 2, and we solve for y: y = 2/(x^2)

OpenStudy (mathstudent55):

Now we go back to our expression for the area, and we replace y with 2/(x^2): A = x^2 + 4xy A = x^2 + 4x[2/(x^2)] A = x^2 + 8x/x^2 A = x^2 + 8/x

OpenStudy (anonymous):

wow you actually made that make sense!

OpenStudy (mathstudent55):

That's great. Go over it step by step. It's important to learn the method.

OpenStudy (anonymous):

Oh I will

OpenStudy (mathstudent55):

Great.